OpenStudy (anonymous):

Lenny is planning to cut down a pine tree. He has parked his truck 90 feet away from the base of the tree and wants to make sure that the tree will not hit his truck when it falls. The angle of elevation from the base of the truck to the top of the tree is 50°. How tall is the tree to the nearest foot? Should Lenny move his truck?

OpenStudy (anonymous):

Let x be the height of the tree. tan 50 = x/90 1.19 = x/90 107 = x The tree is 107 feet tall. He should move his truck at least 17 feet further away from the tree.